OpenStudy (anonymous):

Find all complex solutions of 3x^2+x+4=0 I have -4 did I do this right?

OpenStudy (wolf1728):

a = 3 b = 1 c = 4 x= -1 +- sqrt(1 - 4*3*4) / (2*3) x = -1 +- sqrt (-47) / 6 x1 = (-1/6) + i* (6.8556546004 / 6) x2 = (-1/6) - i*(6.8556546004 /6)

OpenStudy (wolf1728):

So, both solutions are complex and so you weren't right with an answer of -4

OpenStudy (anonymous):

Thank you!

OpenStudy (wolf1728):

And here's a calculator to see if you have done the work correctly: http://www.1728.org/quadratc.htm